2. Assume you are tr compare Project Alpho to Project Beta. Both projects hove the some discount rote. Project Alpha generates infinike after tox cash flows which dlo not grow the 0M one year from no (T 1) ond wil cost $600M today (T-o). Project Beta first cash flow is $8 Penerates infinite ofter tax cash flows which do not grow; the first cosh flow is $40M one year from now t about the crossover rate for Project Alpho and T-1) and will cost $300M soday (T hich s oieen ,os belos me crossover rate o, las Project Alpho projects wos below the crossover rate of 44%, Project Alpho rate of 44% Project Alpho rote of 13,3% Project Alpho projechs wos belox the cromover rate of 20.0% Project Alpha e discount rate for both would have a better b b. If the approp would have a better If the appropriate discount would have o d. If the would have a better on your credit card for 1.0 yeor d compony currently charges you an APR of 200%. If you make efforts to improve credit score by 100 points, your credit cord c compony wa odist your APR by 40%. As me result of you efforts, how much less interest would you pay over 1.0 year o. $1,386 b. $2,000 c. $2,356 d. $2,443 e. $3,412 Issued $400M of long term debt; as port of this inter y lean, the company agreed to pay creditors l 5.0% per year; this implies the current portion of long term debe is SO in 201 7/2018 S400M in sales, $260M in operating earnings, ond $160 in net income .$500M Gross PPE, $450 Net PPE Working Copital of $430, excluding morketoble securiaie Investors were pleased with the results; Company Xs share price closed at $28 per share $4BOM in sales, $300M in operating earnings, and $200 in net income . The company poid o dividend of 30 cents per shore At the end of 2018, Company X issued financial stotements. A complete summary is below $600M Gross PPE, $490 Net PP .Working Copital of $460, excluding marketable securities - The paid o dividend of 40 cents per shore investors were pleased with the results, Compony Xs share price closed at $33 per shore Throughout 2017 and 2018 . The compony did not issue or repurchase ony shores of stock other thon during the PO The compony did not issue or retire ony long term debt other than during the leitial debt offering
. At the end of 2017, Company Xs P/B multiple was closest to 7.8 8.0 8.3 9.3 10.0 5. At the end of 2017 . $90 . $50 $40 . $30 . $20 16. Ai the end of a. $430 b. $420 c $360 d. $320 e. $310 17.At the end a. 20% b. 2396 c. 25% d, 27% e, 33% 18. Do not include marketable securities or the current portion of long term debt when colculating working In 2018, Company Xs Cash Flow from Operations was closest to a. $180 b. $220 c $230 d. $280 e. $290 te question 18 and assume that CFO was $250. Calculate fixed copital investment as the year on year change in Gross PPE. In 2018, Company Xs Free cash Flow to Firm was closest to a. $110 b. $150 c. $170 d. $190 e.$210 20. In 2018, Compony Xs Return on Equity, where ROE-NI/EOY SE, is closest to a, 20% b. 23% c, 25% d.27% e. 33%
Answer #1

Cross over rate is the rate at which NPV of two projects are equal. and

NPV = PV of Inflows - PV of Outflows

Here both the projects forms an infinite cash inflow. So PV of Inflows can be calculated as

t1 cash flow / Discount rate

So Cross over rate (say Ke) is calculated as follows

NPV of project alpha = NPV of project beta

80M/ke - 600M = 40M/ke - 300M

80M/ke - 40M/ke = - 300M + 600M

(80M-40M)/Ke = 300M

40M/Ke = 300M

Ke = 40M/300M

Cross over rate= 13.3%

By using trial and error method;

a. If the appropriate discount rate for both the projects was below the crossover rate of 13.3%, Project Alpha would have a better NPV than Project Beta.

Note: As per HOMEWORKLIB RULES, when multiple questions are posted, we are mandated do only the first question. So please post the questions individually.